# GPU-Accelerated Gravitational Lensing &\&& Dynamical (GLaD) Modeling for Cosmology and Galaxies

### v1

@@
Time-delay distance measurements from strongly lensed quasars provide a robust and independent method for determining the Hubble constant ($H_{0}$). This approach offers a crucial cross-check against $H_{0}$ measurements obtained from the standard distance ladder in the late universe and the cosmic microwave background in the early universe. However, the mass-sheet degeneracy in strong lensing models may introduce significant systematic uncertainty, limiting the precision of $H_{0}$ estimates. Dynamical modeling highly complements strong lensing to break the mass-sheet degeneracy, as both methods model the mass distribution of galaxies but rely on different sets of observational constraints. In this study, we develop a methodology and software framework for efficient joint modeling of stellar kinematic and lensing data. Using simulated lensing and kinematic data of the lensed quasar system RXJ1131--1131 as a test case, we demonstrate that approximately 4% precision on $H_{0}$ is achievable with high-quality and signal-to-noise data. Through extensive modeling, we examine the impact of the presence of a supermassive black hole in the lens galaxy and potential systematic biases in kinematic data on $H_{0}$ measurements. Our results demonstrate that either using a prior range for black hole mass and orbital anisotropy, as motivated by studies of nearby galaxies, or excluding the central bins in the kinematic data, can both effectively mitigate potential biases on $H_{0}$ induced by the black hole. By testing on mock kinematic data with values that are systematically biased, we emphasize the importance of using kinematic data with systematic errors under sub-percent control, which is currently achievable. Additionally, we leverage GPU parallelization to accelerate Bayesian inference, reducing a previously month-long process by an order of magnitude. This pipeline offers significant potential for advancing cosmological and galaxy evolution studies with large datasets.
@@ 1 Introduction
The Hubble constant, $H_{0}$, sets the local expansion rate of the universe and plays a crucial role in understanding its age and size. Previous studies have reported a significant $5\sigma$ tension between $H_{0}$ measurements from the cosmic microwave background (CMB), which gives $H_{0}=67.4\pm 0.5\,\rm km\,s^{-1}\,Mpc^{-1}$ (e.g., Planck Collaboration et al. 2020), and local distance indicators, such as supernovae (SNe) and Cepheid variables, which yield $H_{0}=73.0\pm 1.0\,\rm km\,s^{-1}\,Mpc^{-1}$ (e.g., Riess et al. 2022). However, recent measurements from the Chicago-Carnegie Hubble Program (e.g., Freedman et al. 2024), which are also based on the SN distance ladder, show no significant tension with the CMB, leaving the true discrepancy uncertain. Riess et al. (2024) highlighted that the $H_{0}$ measurement in Freedman et al. (2024) was based on a subsample selection. Whether the tension is real, or merely a result of systematic uncertainties that were not known and not incorporated in the measurements, remains a topic of debate (Efstathiou & Gratton 2020; Abdalla et al. 2022; Yeung & Chu 2022; Freedman & Madore 2023), but if confirmed, it would indicate the need for new physics beyond the standard cosmological model.
@@
Time-delay cosmography offers a distinct approach, separate from the previously mentioned methods, to measure $H_{0}$ by analyzing the brightness variations of sources like quasars or supernovae. It constrains cosmological parameters by measuring the time delay between multiple lensed images of the source (Refsdal 1964; Meylan et al. 2006; Treu & Marshall 2016; Treu et al. 2022; Treu & Shajib 2023; Birrer et al. 2024; Oguri 2019; Liao et al. 2022; Suyu et al. 2024). By determining the time-delay distance to the lens system, it is possible to infer the value of $H_{0}$. However, this approach is affected by the mass-sheet degeneracy (MSD) in strong lensing (SL) (e.g., Falco et al. 1985; Gorenstein et al. 1988; Birrer et al. 2016; Chen et al. 2021a). We categorize MSD into two types: external and internal. Both can potentially bias estimates of $H_{0}$. The external MSD, which arises from line-of-sight (LoS) effects, can be controlled by studying the environments of the lens galaxies (e.g., Wells et al. 2024). The internal MSD arises from the unknown radial profile of the lens galaxies’ mass distribution (e.g., Schneider & Sluse 2013). This degeneracy allows for equally well-fitting models of the observed lensing data while introducing a linear bias in the inferred value of $H_{0}$.
@@
A common strategy to address the internal MSD is to incorporate independent datasets, such as kinematic or weak lensing data (e.g., Treu & Koopmans 2002; Shajib et al. 2020; Birrer et al. 2020; Birrer & Treu 2021; Yıldırım et al. 2023; Shajib et al. 2023; Khadka et al. 2024). These additional observations help detect changes in the mass density slope induced by the internal MSD in SL at the inner region within the Einstein radius $R_{\rm Ein}$ and the outer region $\sim 8R_{\rm Ein}$ from the lens galaxy’s centroid, allowing for a more robust constraint on the mass distribution and, consequently, on $H_{0}$.
@@
The H0LiCOW collaboration reported an $H_{0}$ measurement with $2.4\%$ precision by combining six lensed quasar systems (Wong et al. 2020). These analyses tested two specific mass models, the composite model (baryonic component + dark matter) and the elliptical power-law model, while performing lens modeling without explicitly accounting for the internal MSD. Intuitively, explicitly modeling the internal MSD makes the adopted mass model more flexible, allowing for a broader range of mass distributions. High-resolution spatial kinematics can help distinguish between these more flexible models. However, H0LiCOW used slit kinematics, which primarily served to validate the best-fit mass models rather than to differentiate between them, as slit kinematics alone is insufficient to break the MSD and measure distances with a few-percent uncertainty on an individual lens basis. If mass model assumptions are relaxed and an internal mass sheet—maximally degenerate with $H_{0}$—is incorporated, the precision of the $H_{0}$ constraint from the six lensed quasar systems degrades to $5\%$ or $8\%$, depending on whether external priors from non-time-delay lenses are used for orbital anisotropy (Birrer et al. 2020). The TDCOSMO collaboration continues to investigate potential degeneracies and biases in the measurement of $H_{0}$ caused by the internal MSD in lens modeling (e.g., Millon et al. 2020; Birrer & Treu 2021; Chen et al. 2021a; Van de Vyvere et al. 2022; Gomer et al. 2022), previously studied by the H0LiCOW collaboration. As part of TDCOSMO, Shajib et al. (2023) conducted a joint modeling analysis to explicitly break the internal MSD using spatially resolved kinematics from KCWI (an integral field spectrograph at Keck (Morrissey et al. 2018)). Their study yielded a value of $H_{0}=77.1_{-7.1}^{+7.3}\leavevmode\nobreak\ \rm km\leavevmode\nobreak\ s^{-1},Mpc^{-1}$, achieving a precision of approximately $9.5\%$, from a single time-delay lens system. This analysis was constrained by the kinematic resolution of KCWI. [^1] The kinematic data exhibit a signal-to-noise ratio of $23\leavevmode\nobreak\ \AA^{-1}$ in the rest-frame wavelength range $3985-4085\leavevmode\nobreak\ \AA$ across 41 bins, with a seeing effect of $0.96\arcsec$ in full width at half maximum (FWHM). The diffraction-limited resolution of JWST will offer significantly greater precision, further enhancing kinematic constraints.
@@
The TDCOSMO collaboration aims to constrain $H_{0}$ to within 2% by combining spatially resolved kinematics data obtained from the JWST NIRSpec IFU for seven gravitational lenses. In order to achieve this level of precision and accuracy, extensive tests have been conducted, including examining the impact of the field of view (FoV) on kinematics, comparing different mass models, such as the composite and power-law models and evaluating various dark matter profiles, including the standard NFW profile and its generalized form (e.g., Yıldırım et al. 2020, 2023). Additionally, the influence of the deprojected 3D shape of lens galaxies has been investigated (Shajib et al. 2023; Huang et al. 2025). Exploring these effects requires substantial computational resources, making joint modeling highly demanding. For a single lensed-quasar system such as RXJ1131, Bayesian inference using the Markov chain Monte Carlo (MCMC) method takes a month to complete using traditional CPU-based methods.
@@ 2 Overview of the lens and dynamical modeling
In Sect. 2.1, we provide a brief overview of the SL formalism in the context of time-delay cosmography. In Sect. 2.2, we introduce the MSD, a major source of systematic uncertainty in SL modeling that limits the precision of $H_{0}$ measurements. The internal MSD arises from the unknown size and brightness of source galaxies, as well as the uncertain mass distribution of lens galaxies. These uncertainties impact the measurement of the time-delay distance $D_{\Delta\rm t}$, which is directly proportional to $H_{0}^{-1}$. Similarly, the external MSD, caused by unknown mass distributions along the LoS, introduces an additional uncertainty in $D_{\Delta\rm t}$ measurement, as discussed in Sect. 2.3. In this section, we also demonstrate that the external MSD does not affect dynamical modeling at the galaxy scale, where both lensing and kinematic data are available. In Sect. 2.4, we provide a brief overview of stellar dynamical modeling, assuming an axisymmetric mass distribution and employing the Jeans Anisotropic Modeling (JAM) approach (Cappellari 2008, 2020).
@@ 2.2 Internal mass sheet degeneracy
leaving observables such as image positions and the morphology of lensed image invariant under this transformation, which is known as the internal MSD (Falco et al. 1985). In other words, suppose we model the surface mass density of the lens galaxy as $\kappa_{\rm gal}$ (e.g., using a power-law profile), then $\kappa_{\rm int}$ that accounts for the internal mass sheet would fit lensed image positions and morphology equally well. This transformation propagates to the lens potential via Poisson’s equation:
@@
$D_{\Delta\rm t,int}=\frac{D_{\Delta\rm t,gal}}{\lambda_{\rm int}}\propto\frac{1}{\lambda_{\rm int}H_{0}},$ (9)
where $D_{\Delta\rm t,gal}$ is associated with $\kappa_{\rm gal}$ and $D_{\Delta\rm t,int}$ with $\kappa_{\rm int}$. The internal MSD alters the mass density slope of lens galaxies. This occurs because, aside from the renormalization factor $\lambda_{\rm int}$ in the second term of Eq. 6, the first term results in the addition of a constant sheet to the initial $\kappa_{\rm gal}$. Therefore, if the intrinsic radial profile of the mass distribution in lens galaxies were known, the internal MSD would cease to be a degeneracy. However, in practice, the underlying mass distribution may not be known to sufficient precision, making the class of mass models $\kappa_{\rm int}$ indistinguishable from $\kappa_{\rm gal}$ when relying solely on lensing data. In time-delay cosmography, this means that $D_{\Delta\rm t,int}$ yields a linearly scaled $\lambda_{\rm int}H_{0}$.
@@
2.3 External mass sheet degeneracy
Unlike internal MSD, which has relatively strong effects on small scales, such as altering mass density slopes of lens galaxies, external MSD merely performs the renormalisation of the underlying mass convergence distribution. We use a class of $\kappa_{\rm int}$ to represent the mass distributions of lens galaxies, as they are all viable choices until distinguished by kinematic data. In the external MSD regime, $\kappa_{\rm int}$ scales as:
$\kappa_{\rm int,ext}=(1-\kappa_{\rm ext})\kappa_{\rm int}+\kappa_{\rm ext}$ (10)
@@ 2.3 External mass sheet degeneracy
$D_{\Delta\rm t,int,ext}=\frac{D_{\Delta\rm t,int}}{(1-\kappa_{\rm ext})}.$ (11)
The external convergence $\kappa_{\rm ext}$ can be estimated by examining the lens environment using photometric and spectroscopic surveys, as well as through ray-tracing methods in cosmological simulations (e.g., Suyu et al. 2010; Greene et al. 2013; Suyu et al. 2014; Rusu et al. 2017). We investigate whether the renormalization factor $(1-\kappa_{\rm ext})$ from the external MSD affects the dynamical modeling. We derive the 2D surface mass density $\Sigma_{\rm int,ext}$ as
$\Sigma_{\rm int,ext}=\Sigma_{\rm crit}\kappa_{\rm int,ext}=\Sigma_{\rm crit}\left[(1-\kappa_{\rm ext})\kappa_{\rm int}+\kappa_{\rm ext}\right],$ (12)
where $\Sigma_{\rm crit}$ is the critical density. In the framework of external MSD, we express $\Sigma_{\rm crit}$ in terms of $D_{\rm\Delta t,int,ext}$ 444 $D_{\Delta\rm t,int,ext}$ represents the actual distance, i,e, the distance that can be directly compared to the predictions from cosmological models. as
$\Sigma_{\rm crit}=\frac{c^{2}}{4\pi G}\frac{D_{\rm\Delta t,int,ext}}{(1+z_{\rm d})D_{\rm d}^{2}},$ (13)
where $D_{\rm d}$ 555 The value of $D_{\rm d}$ remains also unchanged by internal MSD as it is exclusively derived from the dynamical modeling. remains fully invariant under external MSD (Jee et al. 2015):
$D_{\rm d,ext}=D_{\rm d}.$ (14)
By substituting Eqs. 13 and 14 into Eq. 12, we find that the factor $(1-\kappa_{\rm ext})$ cancels out in the first term of Eq. 12. As a result, the 2D surface mass density $\Sigma_{\rm int,ext}$ simplifies to
$\Sigma_{\rm int,ext}=\frac{c^{2}}{4\pi G}\frac{1}{(1+z_{\rm d})D_{\rm d}^{2}}\left[D_{\rm\Delta t,int}\kappa_{\rm int}+D_{\rm\Delta t,int,ext}\kappa_{\rm ext}\right].$ (15)
In the lensing and dynamical modeling, we focus on modeling the first term in Eq. 15, which remains unaffected by $(1-\kappa_{\rm ext})$. The second term in the Eq. 15 is essentially a constant accounting for all the perturbations along LoS that do not affect the dynamics of the lens galaxy. As a result, constraining the internal MSD parameter $\lambda_{\rm int}$ is independent of the external convergence.

### v2

@@
Time-delay distance measurements from strongly lensed quasars provide a robust and independent method for determining the Hubble constant ($H_{0}$). This approach offers a crucial cross-check against $H_{0}$ measurements obtained from the standard distance ladder in the late universe and the cosmic microwave background in the early universe. However, the mass-sheet degeneracy in strong lensing models may introduce significant systematic uncertainty, limiting the precision of $H_{0}$ estimates. Dynamical modeling highly complements strong lensing to break the mass-sheet degeneracy, as both methods model the mass distribution of galaxies but rely on different sets of observational constraints. In this study, we develop a methodology and software framework for efficient joint modeling of stellar kinematic and lensing data. Using simulated lensing and kinematic data of the lensed quasar system RXJ1131--1131 as a test case, we demonstrate that approximately 4% precision on $H_{0}$ is achievable with high-quality and signal-to-noise data. Through extensive modeling, we examine the impact of the presence of a supermassive BH in the lens galaxy and potential systematic biases in kinematic data on $H_{0}$ measurements. Our results demonstrate that either using a prior range for BH mass and orbital anisotropy, as motivated by studies of nearby galaxies, or excluding the central bins in the kinematic data, can both effectively mitigate potential biases on $H_{0}$ induced by the BH. By testing on mock kinematic data with values that are systematically biased, we emphasize the importance of using kinematic data with systematic errors under sub-percent control, which is currently achievable. Additionally, we leverage GPU parallelization to accelerate Bayesian inference, reducing a previously month-long process by an order of magnitude. This pipeline offers significant potential for advancing cosmological and galaxy evolution studies with large datasets.
@@ 1 Introduction
The Hubble constant, $H_{0}$, sets the local expansion rate of the universe and plays a crucial role in understanding its age and size. Previous studies have reported a significant $5\sigma$ tension between $H_{0}$ measurements from the cosmic microwave background (CMB), which gives $H_{0}=67.4\pm 0.5\,\rm km\,s^{-1}\,Mpc^{-1}$ (e.g., Planck Collaboration et al. 2020), and local distance indicators, such as supernovae (SNe) and Cepheid variables, which yield $H_{0}=73.0\pm 1.0\,\rm km\,s^{-1}\,Mpc^{-1}$ (e.g., Riess et al. 2022). However, recent measurements from the Chicago-Carnegie Hubble Program (e.g., Freedman et al. 2025), which are also based on the SN distance ladder, show no significant tension with the CMB, leaving the true discrepancy uncertain. Riess et al. (2024) highlighted that the $H_{0}$ measurement in Freedman et al. (2025) was based on a subsample selection. Whether the tension is real, or merely a result of systematic uncertainties that were not known and not incorporated in the measurements, remains a topic of debate (Efstathiou & Gratton 2020; Abdalla et al. 2022; Yeung & Chu 2022; Freedman & Madore 2023), but if confirmed, it would indicate the need for new physics beyond the standard cosmological model.
@@
Time-delay cosmography offers a distinct approach, separate from the previously mentioned methods, to measure $H_{0}$ by analyzing the brightness variations of sources like quasars or supernovae. It constrains cosmological parameters by measuring the time delay between multiple lensed images of the source (Refsdal 1964; Meylan et al. 2006; Treu & Marshall 2016; Treu et al. 2022; Treu & Shajib 2024; Birrer et al. 2024; Oguri 2019; Liao et al. 2022; Suyu et al. 2024). By determining the time-delay distance to the lens system, it is possible to infer the value of $H_{0}$. However, this approach is affected by the mass-sheet degeneracy (MSD) in strong lensing (SL) (e.g., Falco et al. 1985; Gorenstein et al. 1988; Birrer et al. 2016; Chen et al. 2021a). We categorize MSD into two types: external and internal. Both can potentially bias estimates of $H_{0}$. The external MSD, which arises from line-of-sight (LoS) effects, can be controlled by studying the environments of the lens galaxies (e.g., Wells et al. 2024). The internal MSD arises from the unknown radial profile of the lens galaxies’ mass distribution (e.g., Schneider & Sluse 2013). This degeneracy allows for equally well-fitting models of the observed lensing data while introducing a linear bias in the inferred value of $H_{0}$.
@@
A common strategy to address the internal MSD is to incorporate independent datasets, such as kinematic or weak lensing data (e.g., Treu & Koopmans 2002; Shajib et al. 2020; Birrer et al. 2020; Birrer & Treu 2021; Yıldırım et al. 2023; Shajib et al. 2023; Khadka et al. 2024). These additional observations help detect changes in the mass density slope induced by the internal MSD in SL, both in the inner regions covered by the kinematic measurements—depending on the field of view (FoV) and the signal-to-noise ratio—and in the outer regions covered by weak lensing, up to $\sim 8R_{\rm Ein}$ from the lens galaxy’s centroid. This enables a more robust constraint on the mass distribution and, consequently, on $H_{0}$.
@@
The H0LiCOW collaboration reported an $H_{0}$ measurement with $2.4\%$ precision by combining six lensed quasar systems (Wong et al. 2020). These analyses tested two specific mass models, the composite model (baryonic component + dark matter) and the elliptical power-law model, while performing lens modeling without explicitly accounting for the internal MSD. Intuitively, explicitly modeling the internal MSD makes the adopted mass model more flexible, allowing for a broader range of mass distributions. High-resolution spatial kinematics can help distinguish between these more flexible models. However, H0LiCOW used slit kinematics, which primarily served to validate the best-fit mass models rather than to differentiate between them, as slit kinematics alone is insufficient to break the MSD and measure distances with a few-percent uncertainty on an individual lens basis. If mass model assumptions are relaxed and an internal mass sheet—maximally degenerate with $H_{0}$—is incorporated, the precision of the $H_{0}$ constraint from the six lensed quasar systems degrades to $5\%$ or $8\%$, depending on whether external priors from non-time-delay lenses are used for orbital anisotropy (Birrer et al. 2020). The TDCOSMO collaboration continues to investigate potential degeneracies and biases in the measurement of $H_{0}$ caused by the internal MSD in lens modeling (e.g., Millon et al. 2020; Birrer & Treu 2021; Chen et al. 2021a; Van de Vyvere et al. 2022; Gomer et al. 2022), previously studied by the H0LiCOW collaboration. As part of TDCOSMO, Shajib et al. (2023) conducted a joint modeling analysis to explicitly break the internal MSD using spatially resolved kinematics from KCWI (an integral field spectrograph at Keck Observatory (Morrissey et al. 2018)). Their study yielded a value of $H_{0}=77.1_{-7.1}^{+7.3}\leavevmode\nobreak\ \rm km\leavevmode\nobreak\ s^{-1},Mpc^{-1}$, achieving a precision of approximately $9.5\%$, from a single time-delay lens system. This analysis was constrained by the kinematic resolution of KCWI. [^1] The kinematic data exhibit a signal-to-noise ratio of $23\leavevmode\nobreak\ \AA^{-1}$ in the rest-frame wavelength range $3985-4085\leavevmode\nobreak\ \AA$ across 41 bins, with a seeing effect of $0.96\arcsec$ in full width at half maximum (FWHM). The diffraction-limited resolution of JWST will offer significantly greater precision, further enhancing kinematic constraints.
@@
The TDCOSMO collaboration aims to constrain $H_{0}$ to within 2% by combining spatially resolved kinematics data obtained from the JWST NIRSpec IFU for seven gravitational lenses. In order to achieve this level of precision and accuracy, extensive tests have been conducted, including examining the impact of the FoV on kinematics, comparing different mass models, such as the composite and power-law models and evaluating various dark matter profiles, including the standard NFW profile and its generalized form (e.g., Yıldırım et al. 2020, 2023). Additionally, the influence of the deprojected 3D shape of lens galaxies has been investigated (Shajib et al. 2023; Huang et al. 2025). Exploring these effects requires substantial computational resources, making joint modeling highly demanding. For a single lensed-quasar system such as RXJ1131, Bayesian inference using the Markov chain Monte Carlo (MCMC) method takes a month to complete using traditional CPU-based methods.
@@ 2 Overview of the lens and dynamical modeling
In Sect. 2.1, we provide a brief overview of the SL formalism in the context of time-delay cosmography. In Sect. 2.2, we introduce the MSD, a major source of systematic uncertainty in SL modeling that limits the precision of $H_{0}$ measurements. The internal MSD arises from the unknown size and brightness of source galaxies, as well as the uncertain mass distribution of lens galaxies. These uncertainties impact the measurement of the time-delay distance $D_{\Delta\rm t}$, which is directly proportional to $H_{0}^{-1}$. Similarly, the external MSD, caused by unknown mass distributions along the LoS, introduces an additional uncertainty in $D_{\Delta\rm t}$ measurement, as discussed in Sect. 2.3. In this section, we also show that when using joint modeling with a non-fixed cosmological model, the external MSD does not impact dynamical modeling. In Sect. 2.4, we provide a brief overview of stellar dynamical modeling, assuming an axisymmetric mass distribution and employing the Jeans Anisotropic Modeling (JAM) approach (Cappellari 2008, 2020).
@@ 2.2 Internal mass sheet degeneracy
leaving observables such as image positions and the morphology of lensed images invariant under this transformation, which is known as the internal MSD (Falco et al. 1985). In other words, suppose we model the surface mass density of the lens galaxy as $\kappa_{\rm gal}$ (e.g., using a power-law profile), then $\kappa_{\rm int}$ that accounts for the internal mass sheet would fit lensed image positions and morphology equally well. This transformation propagates to the lens potential via Poisson’s equation:
@@
$D_{\Delta\rm t,int}$ $=\frac{D_{\Delta\rm t,gal}}{\lambda_{\rm int}}=(1+z_{\rm d})\frac{D_{\rm d,int}D_{\rm s,int}}{D_{\rm ds,int}}$ (9)
$\propto\frac{1}{\lambda_{\rm int}H_{0}}$
where $D_{\Delta\rm t,gal}$ is associated with $\kappa_{\rm gal}$ and $D_{\Delta\rm t,int}$ with $\kappa_{\rm int}$. The distances $D_{\rm s,int}$, $D_{\rm ds,int}$, and $D_{\rm d,int}$ are considered in the context of the internal MSD associated with $D_{\rm\Delta t,int}$. Note that the distance to lens galaxy is influenced by internal MSD, such that,
$D_{\rm d,gal}\neq D_{\rm d,int}.$ (10)
However, the change of $D_{\rm d}$ induced by internal MSD is minor under $1\%$ shift, given the single aperture kinematics (Chen et al. 2021a, see Sect. 4.2). With 2D resolved kinematics, the impact on $D_{\rm d,int}$ could be more significant.
The internal MSD alters the mass density slope of lens galaxies. This occurs because, aside from the renormalization factor $\lambda_{\rm int}$ in the second term of Eq. 6, the first term results in the addition of a constant sheet to the initial $\kappa_{\rm gal}$. Therefore, if the intrinsic radial profile of the mass distribution in lens galaxies were known, the internal MSD would cease to be a degeneracy. However, in practice, the underlying mass distribution may not be known to sufficient precision, making the class of mass models $\kappa_{\rm int}$ indistinguishable from $\kappa_{\rm gal}$ when relying solely on lensing data. In time-delay cosmography, this means that $D_{\Delta\rm t,int}$ yields a linearly scaled $\lambda_{\rm int}H_{0}$.
@@
2.3 Internal and External mass sheet degeneracy
Unlike internal MSD, which has relatively strong effects on small scales, such as altering mass density slopes of lens galaxies, external MSD merely performs the renormalization of the underlying mass convergence distribution. We use a class of $\kappa_{\rm int}$ to represent the mass distributions of lens galaxies, as they are all viable choices until distinguished by kinematic data. In the external MSD regime, $\kappa_{\rm int}$ scales as:
$\kappa=(1-\kappa_{\rm ext})\kappa_{\rm int}+\kappa_{\rm ext}$ (11)
@@ 2.3 Internal and External mass sheet degeneracy
$D_{\Delta\rm t}=\frac{D_{\Delta\rm t,int}}{(1-\kappa_{\rm ext})}.$ (12)
The inferred $D_{\Delta\rm t}$ is a distance that can be compared with the distance calculated from the cosmological models and test the cosmology, whereas $D_{\Delta\rm t,int}$ cannot be used to test assumed cosmological models directly because it has not been corrected for the external convergence. The external convergence $\kappa_{\rm ext}$ can be estimated by examining the lens environment using photometric and spectroscopic surveys, as well as through ray-tracing methods in cosmological simulations (e.g., Suyu et al. 2010; Greene et al. 2013; Suyu et al. 2014; Rusu et al. 2017). We investigate whether the renormalization factor $(1-\kappa_{\rm ext})$ from the external MSD affects the dynamical modeling. We derive the 2D surface mass density $\Sigma$ as
$\Sigma=\Sigma_{\rm crit}\kappa=\Sigma_{\rm crit}\left[(1-\kappa_{\rm ext})\kappa_{\rm int}+\kappa_{\rm ext}\right],$ (13)
where $\Sigma_{\rm crit}$ is the critical density. In the framework of internal and external MSD, we express $\Sigma_{\rm crit}$ in terms of $D_{\rm\Delta t}$ 444 $D_{\Delta\rm t,int}$ represents the actual distance, i,e, the distance that can be directly compared to the predictions from cosmological models. as
$\Sigma_{\rm crit}=\frac{c^{2}}{4\pi G}\frac{D_{\rm\Delta t}}{(1+z_{\rm d})D_{\rm d}^{2}},$ (14)
where $D_{\rm d}$ remains fully invariant under external MSD when modeling is performed without fixing the cosmology, using lens dynamics and time-delay cosmography jointly (Jee et al. 2015; Chen et al. 2021a):
$D_{\rm d}=D_{\rm d,int}.$ (15)
By substituting Eqs. 14 and 15 into Eq. 13, we find that the factor $(1-\kappa_{\rm ext})$ cancels out in the first term of Eq. 16. As a result, the 2D surface mass density $\Sigma$ is written as
$\Sigma$ $=\frac{c^{2}}{4\pi G}\frac{D_{\Delta\rm t}}{(1+z_{\rm d})D_{\rm d}^{2}}\left[(1-\kappa_{\rm ext})\kappa_{\rm int}+\kappa_{\rm ext}\right]$
$=\frac{c^{2}}{4\pi G}\frac{D_{\rm\Delta t,\rm int}}{(1+z_{\rm d})(1-\kappa_{\rm ext})D_{\rm d,int}^{2}}\left[(1-\kappa_{\rm ext})\kappa_{\rm int}+\kappa_{\rm ext}\right]$
$=\frac{c^{2}}{4\pi G}\frac{1}{(1+z_{\rm d})D_{\rm d,int}^{2}}\left[D_{\Delta\rm t,\rm int}\kappa_{\rm int}+D_{\Delta\rm t}\kappa_{\rm ext}\right]$
$=\Sigma_{\rm int}+\Sigma_{\rm ext}$ (16)
In the lensing and dynamical modeling, we focus on modeling $\Sigma_{\rm int}$ in Eq. 16. We sample the distance $D_{\Delta\rm t,\rm int}$ in the joint modeling, rather than $D_{\Delta\rm t}$, to ensure that the dynamical modeling remains unaffected by the external convergence $\kappa_{\rm ext}$. Sampling $D_{\Delta\rm t}$ directly will introduce a scaling factor of $(1-\kappa_{\rm ext})$ into the dynamical analysis. The second term $\Sigma_{\rm ext}$ in Eq. 16 is essentially a constant accounting for all the perturbations along LoS that do not affect the dynamics of the lens galaxy.